# FlexRT - A fast and flexible cosmological radiative transfer code for reionization studies I: Code validation

### v1

@@ 5.1 Iso-thermal Stromgren Sphere (Test #1)
Figure 4 shows the angle-averaged HI fraction ($x_{\rm HI}$) profile as a function of distance from the source. The left and right panels show results at $\Delta t=30$ and $500$ Myr, respectively. Note that in Generalized Opacity mode, the HI fraction is not calculated explicitly, and so we must invert Eq. 5.5 to get it. The two modes agree well with each other at $t=30$ Myr, when the I-front is sharpest. At $t=500$ Myr, when the I-front has nearly stopped, a larger difference emerges. The transition from highly ionized to neutral is wider in Standard mode and extends to a larger distance from the source. Behind the I-front, the profile of residual HI is is good agreement. Standard mode is in tight agreement with several CRTCP modes, particularly C2-Ray. This makes sense because standard mode and C2-Ray have very similar ionization solvers. Once again, Generalized Opacity mode agrees best with IFT.
@@
Figure 5 shows slices of the HI fraction through the center of the box at $t=30$ and $500$ Myr (top and bottom rows, respectively)111111 We only show one corner of the HII region, since the CRTCP tests do not simulate the entire sphere.. In the 1st and 2nd column, we compare Standard mode to C2-Ray, since these use similar methods to solve for the ionization state of the gas. The 3rd and 4th columns compare Generalized Opacity mode to IFT, since these both assume sharp boundaries between ionized and neutral gas. We find excellent visual agreement between FlexRT and C2-Ray at both times. Generalized Opacity mode displays a much sharper boundary between ionized and neutral gas, especially at $\Delta t=30$ Myr, reflecting its sharp I-front treatment. In IFT, there is a similarly sharp jump between highly ionized ($x_{\rm HI}<0.01$) gas and highly neutral gas, although $x_{\rm HI}$ does not jump immediately to unity as it does in Generalized Opacity mode. There is also some noise noticeable in Generalized Opacity mode. This arises from numerical artifacts associated with the HealPix treatment of the radiation field. In Generalized Opacity mode, $n_{\rm HI}$ is estimated at each time step directly from the radiation field (Eq. 5.7), so these effects manifest in $x_{\rm HI}$ more clearly than they do in Standard mode, where $x_{\rm HI}$ is time-integrated (Eq. 4.6).
@@ 6 Validity of the Moving-Screen I-front method
The moving-screen I-front approximation assumes that the I-front is entirely unresolved - namely, the typical width of an I-front ($\Delta x_{\rm IF}$) is much smaller than $\Delta x_{\rm cell}$. To demonstrate this, we have run a cosmological simulation in a $200$ $h^{-1}$pc box with a cosmological distribution of halos acting as ionizing sources. The cells in the box have $\Delta x_{\rm cell}=1$ $h^{-1}$pc, considerably larger than the widths of typical cosmological I-fronts ($\Delta x_{\rm IF}\sim 10-100$ $h^{-1}$kpc). The details of this simulation are described in the next section, so we will omit them here - save to emphasize that the moving-screen approximation is expected to be valid in this case because of the large RT cells. The top-left and top-right panels of Figure 15 show maps of $x_{\rm HI}$ (in log scale) for this test in Standard and Generalized Opacity modes, respectively. The inset shows a zoom-in to highlight the structure of the I-fronts. In both cases, the boundary between highly ionized (blue) and neutral (red) cells is only one cell thick, showing that the I-front is un-resolved. The two modes agree very well in this limit.
@@
Figure 15: Demonstration of the regime in which the moving-screen I-front approximation used in Generalized Opacity mode is valid. In the top row, we show maps of $x_{\rm HI}$ from a cosmological simulation in a $200$ $h^{-1}$pc with $\Delta x_{\rm cell}=1$ $h^{-1}$pc in both modes. The insets show zoom-ins to highlight the structure of I-fronts. The boundary between neutral (red) and highly ionized (blue/green) cells is one cell thick, indicating that the I-fronts are un-resolved. In this limit, Standard and Generalized Opacity modes produce the same result. In the bottom row, we scale the box volume and ionizing emissivity down by a factor of $100^{3}$, so that $\Delta x_{\rm cell}=10$ $h^{-1}$kpc. In this case, the I-fronts are partially resolved in Standard mode (bottom left), differing from the result of the moving-screen approximation seen in the bottom right. This comparison shows that Generalized Opacity mode should be applied in cases where I-fronts are un-resolved. This turns out to be its intended regime of applicability.
@@
In the bottom panels, we show a contrasting example in which the moving-screen approximation fails. To do this, we artificially scale the volume of our $200$ $h^{-1}$pc box down by a factor of $100^{3}$ so that it becomes a $2$ $h^{-1}$pc box. We also scale the ionizing emissivity of sources down by the same factor so that reionization progresses at the same rate. We then rerun the simulation in the two modes. These runs have $\Delta x_{\rm cell}=10$ $h^{-1}$kpc, such that I-fronts are partially resolved. We see that in the bottom-left panel, the boundary between highly ionized and fully neutral gas is several cells thick, whereas it remains one cell thick on the bottom-right. This comparison illustrates the regime in which the moving-screen approximation employed in Generalized Opacity mode is valid: $\Delta x_{\rm cell}>>\Delta x_{\rm IF}$. We emphasize that Generalized Opacity mode is designed for large-volume simulations of reionization (see e.g. [21, 35]). Standard mode should be used for problems in which the I-fronts are resolved and their detailed structure is important.
@@ 7.2 Test against RadHydro
Figure 20: Top: auto power spectrum of the ionization field $\Delta_{\rm ion}(k)$ for FlexRT and RadHydro (solid black and dotted red), and their cross-correlation (dashed blue). Bottom: their cross-correlation coefficient, $r(k)$ (Eq. 7.1). From left to right, the columns show results at $20\%$, $50\%$, and $80\%$ ionized. At $20\%$ and $50\%$ ionized, the power spectra are almost indistinguishable, with $r(k)$ differing from unity only close to the Nyquist wavenumber ($k_{\rm nyq}=\pi/\Delta x_{\rm cell}\approx 3$ $h$Mpc-1). At $80\%$, the codes are still in good agreement, but RadHydro has slightly more ionization power at intermediate scales. This is probably due to differences in how the codes handle ray merging as the ionized fraction nears unity.
We show the ionized fraction as a function of time in the top panel of Figure 18 for RadHydro (black) and FlexRT (blue), and their ratio in the bottom panel. The solid (dashed) lines denote volume (mass)-weighted averages. The differences between FlexRT and RadHydro are a few percent or less. The slight disagreement may arise from the codes’ different ionization solvers. Specifically, RadHydro assumes the optically thin limit of Eq. 4.1 and does not account for evolution of the neutral fraction during a time step (Eq. 4.2-4.6) when calculating the opacities of cells. This is an important difference for our test setup because the large RT cells ($1$ $h^{-1}$pc) result in large opacities in partially neutral cells ($\tau>>1$) and large time steps ($\Delta t=0.6$ Myr).
@@ 8 Conclusions
  • Using a cosmological reionization simulation, we demonstrated the regime in which the moving-screen I-front approximation in Generalized Opacity mode approaches the results of Standard mode. We found that when the RT cell size, $\Delta x_{\rm cell}$, is much larger than typical I-front widths, ionization maps from the two modes are almost indistinguishable. Standard mode should be used for problems in which the I-fronts are resolved and their detailed structure is important.

### v2

@@ 5.1 Iso-thermal Stromgren Sphere (Test #1)
Figure 4 shows the angle-averaged HI fraction ($x_{\rm HI}$) profile as a function of distance from the source. The left and right panels show results at $\Delta t=30$ and $500$ Myr, respectively. Note that in Generalized Opacity mode, the HI fraction is not calculated explicitly, and so we must invert Eq. 5.5 to get it. The two modes agree well with each other at $t=30$ Myr, when the I-front is sharpest. At $t=500$ Myr, when the I-front has nearly stopped, a larger difference emerges. The transition from highly ionized to neutral is wider in Standard mode and extends to a larger distance from the source. Behind the I-front, the profile of residual HI is in good agreement. Standard mode is in tight agreement with several CRTCP modes, particularly C2-Ray. This makes sense because standard mode and C2-Ray have very similar ionization solvers. Once again, Generalized Opacity mode agrees best with IFT.
@@
Figure 5 shows slices of the HI fraction through the center of the box at $t=30$ and $500$ Myr (top and bottom rows, respectively)111111 We only show one corner of the HII region, since the CRTCP tests do not simulate the entire sphere.. In the 1st and 2nd column, we compare Standard mode to C2-Ray, since these use similar methods to solve for the ionization state of the gas. The 3rd and 4th columns compare Generalized Opacity mode to IFT, since these both assume sharp boundaries between ionized and neutral gas. We find excellent visual agreement between FlexRT and C2-Ray at both times. Generalized Opacity mode displays a much sharper boundary between ionized and neutral gas, especially at $\Delta t=30$ Myr, reflecting its sharp I-front treatment. In IFT, there is a similarly sharp jump between highly ionized ($x_{\rm HI}<0.01$) gas and highly neutral gas, although $x_{\rm HI}$ does not jump immediately to unity as it does in Generalized Opacity mode. There is also some noise noticeable in Generalized Opacity mode, particularly at $\Delta t=500$ Myr. This is a manifestation of the “flickering” effect arising from rotating the HealPix unit vectors associated with the source on each time step. The reason these are visible in Generalized Opacity Mode and not Standard Mode is because, in the former, $\lambda$ is calculated using the instantaneous value of $\Gamma_{\rm HI}$ (Eq. 5.5-Eq. 5.7), such that the neutral fraction effectively reacts instantaneously to changes in incident flux. Note that this is not the case in the sub-grid model described in Refs. [21, 35], which includes some dependence of $\lambda$ on the prior history of $\Gamma_{\rm HI}$. In Standard Mode, $n_{\rm HI}$ is the result of time integration (Eq. 4.6), and is thus much less sensitive to this effect. Fortunately, the tight agreement of Generalized Opacity Mode with the analytic expectation in Figure 3 suggests that this does not significantly affect the recombination rate in this test.
@@ 6 Validity of the Moving-Screen I-front method
The moving-screen I-front approximation assumes that the I-front is entirely unresolved - namely, the typical width of an I-front, $\Delta x_{\rm IF}$, is much smaller than $\Delta x_{\rm RT}$. To demonstrate this, we have run a cosmological simulation in a $200$ $h^{-1}$pc box with a cosmological distribution of halos acting as ionizing sources. The cells in the box have $\Delta x_{\rm RT}=1$ $h^{-1}$pc, considerably larger than the widths of typical cosmological I-fronts ($\Delta x_{\rm IF}\sim 10-100$ $h^{-1}$kpc). The details of this simulation are described in the next section, so we will omit them here - save to emphasize that the moving-screen approximation is expected to be valid in this case because of the large RT cells. The top-left and top-right panels of Figure 15 show maps of $x_{\rm HI}$ (in log scale) for this test in Standard and Generalized Opacity modes, respectively. The inset shows a zoom-in to highlight the structure of the I-fronts. In both cases, the boundary between highly ionized (blue) and neutral (red) cells is only one cell thick, showing that the I-front is un-resolved. The two modes agree very well in this limit.
@@
Figure 15: Demonstration of the regime in which the moving-screen I-front approximation used in Generalized Opacity mode is valid. In the top row, we show maps of $x_{\rm HI}$ from a cosmological simulation in a $200$ $h^{-1}$pc with $\Delta x_{\rm RT}=1$ $h^{-1}$pc in both modes. The insets show zoom-ins to highlight the structure of I-fronts. The boundary between neutral (red) and highly ionized (blue/green) cells is one cell thick, indicating that the I-fronts are un-resolved. In this limit, Standard and Generalized Opacity modes produce the same result. In the bottom row, we scale the box volume and ionizing emissivity down by a factor of $100^{3}$, so that $\Delta x_{\rm RT}=10$ $h^{-1}$kpc. In this case, the I-fronts are partially resolved in Standard mode (bottom left), differing from the result of the moving-screen approximation seen in the bottom right. This comparison shows that Generalized Opacity mode should be applied in cases where I-fronts are un-resolved. This turns out to be its intended regime of applicability. Conversely, the comparison of the case with large $\Delta x_{\rm cell}$ shows that the I-front “smearing” effect discussed in §4.1 does not affect the total recombination rate significantly, at least not for this test.
@@
In the bottom panels, we show a contrasting example in which the moving-screen approximation fails. To do this, we artificially scale the volume of our $200$ $h^{-1}$pc box down by a factor of $100^{3}$ so that it becomes a $2$ $h^{-1}$pc box. We also scale the ionizing emissivity of sources down by the same factor so that reionization progresses at the same rate. We then rerun the simulation in the two modes. These runs have $\Delta x_{\rm RT}=10$ $h^{-1}$kpc, such that I-fronts are partially resolved. We see that in the bottom-left panel, the boundary between highly ionized and fully neutral gas is several cells thick, whereas it remains one cell thick on the bottom-right. This comparison illustrates the regime in which the moving-screen approximation employed in Generalized Opacity mode is valid: $\Delta x_{\rm RT}>>\Delta x_{\rm IF}$. We emphasize that Generalized Opacity mode is designed for large-volume simulations of reionization (see e.g. [21, 35]). Standard mode should be used for problems in which the I-fronts are resolved and their detailed structure is important.
Conversely, our test shows that in the case with $\Delta x_{\rm RT}=1$ $h^{-1}$pc, the growth of ionized regions is largely unaffected by the error in the recombination rate inside I-fronts incurred in Standard Mode (see discussion in S4.1). This suggests that, at least in this test, the recombinations taking place in partially ionized cells do not contribute significantly to the total ionizing budget in the simulation. However, we caution that this result may not hold in the case where the recombination rate is being boosted at the sub-grid level to account for missing small-scale structure. In general, a test like the one shown in this section - comparing the moving-screen method to the Standard Mode procedure - can be used to assess the importance of such recombinations and determine if the treatment in Standard Mode is sufficiently accurate for a given situation.
@@ 7.2 Test against RadHydro
Figure 20: Top: auto power spectrum of the ionization field $\Delta_{\rm ion}(k)$ for FlexRT and RadHydro (solid black and dotted red), and their cross-correlation (dashed blue). Bottom: their cross-correlation coefficient, $r(k)$ (Eq. 7.1). From left to right, the columns show results at $20\%$, $50\%$, and $80\%$ ionized. At $20\%$ and $50\%$ ionized, the power spectra are almost indistinguishable, with $r(k)$ differing from unity only close to the Nyquist wavenumber ($k_{\rm nyq}=\pi/\Delta x_{\rm RT}\approx 3$ $h$Mpc-1). At $80\%$, the codes are still in good agreement, but RadHydro has slightly more ionization power at intermediate scales. This is probably due to differences in how the codes handle ray merging as the ionized fraction nears unity.
We show the ionized fraction as a function of time in the top panel of Figure 18 for RadHydro (black) and FlexRT (blue), and their ratio in the bottom panel. The solid (dashed) lines denote volume (mass)-weighted averages. The differences between FlexRT and RadHydro are a few percent or less. The slight disagreement may arise from the codes’ different ionization solvers. Specifically, RadHydro assumes the optically thin limit of Eq. 4.1 and does not account for evolution of the neutral fraction during a time step (Eq. 4.2-4.6) when calculating the opacities of cells. This is an important difference for our test setup because the large RT cells ($1$ $h^{-1}$pc) result in large opacities in partially neutral cells ($\tau>>1$) and large time steps ($\Delta t=0.6$ Myr). Another, related effect could be that of HealPix structural artifacts, discussed in §3.1. RadHydro does not randomly rotate the HealPix spheres used to cast rays on every time step [^15] Instead, it assigns each source a single set of random rotation angles, which are used throughout the simulation. . This could lead to differences in the recombination rate around isolated sources early in the simulation before ionized regions overlap.
@@ 8 Conclusions
  • Using a cosmological reionization simulation, we demonstrated the regime in which the moving-screen I-front approximation in Generalized Opacity mode approaches the results of Standard mode. We found that when the RT cell size, $\Delta x_{\rm RT}$, is much larger than typical I-front widths, ionization maps from the two modes are almost indistinguishable. Standard mode should be used for problems in which the I-fronts are resolved and their detailed structure is important.